### v1

# Superconducting Cloud Chamber

Abstract
We propose a new particle-trajectory detector composed of Josephson junctions, named the superconducting cloud chamber. By measuring the quantum phase difference, this device can detect charged particles with extremely low kinetic energy, providing a new method for detecting slow-moving particles. It can also be utilized to detect millicharged dark matter particles thermalized with the Earth’s environment within the mass range of $10^{5}\sim 10^{10}\,$eV.
Introduction – Millicharged particles can be an interesting dark matter (DM) candidate Holdom:1985ag ; Feldman:2007wj ; McDermott:2010pa ; Izaguirre:2015eya ; Agrawal:2016quu because millicharge naturally ensures the stability of particle. As long as the millicharge of DM is large enough, it can reach thermal equilibrium with the Earth environment through long-range electromagnetic interactions with standard model (SM) particles Neufeld:2018slx ; Pospelov:2020ktu ; Leane:2022hkk ; Berlin:2023zpn . Consequently, the kinetic energy of DM becomes as tiny as the Earth’s temperature ($\sim 0.2\,$eV), which prevents it from surpassing the energy threshold of DM direct detection experiments and thus evading experimental constraints Pospelov:2020ktu ; Berlin:2023zpn . Additionally, such a strong-interacting millicharged DM is thermally produced in the early Universe with a small abundance Aboubrahim:2021ohe . The cosmological observations are not sensitive to the DM with a small fraction Dubovsky:2003yn ; dePutter:2018xte ; Kovetz:2018zan ; Buen-Abad:2021mvc .
Although many new experiments are proposed to detect millicharged DMs Pospelov:2020ktu ; Kim:2007zzs ; Moore:2014yba ; Afek:2020lek ; Budker:2021quh , there are still many parameter spaces with large mass and small fraction for the millicharged DM that need exploration Budker:2021quh . Moreover, current particle detectors are unable to directly observe extremely low-velocity particles, even those SM charged particles, without acceleration. In this paper, we design a novel superconducting cloud chamber made of RF SQUIDs (radio frequency superconducting quantum interference device 34120 ) to target low-speed charged particles. We use natural units with $c=\hbar=1$ throughout the text.
Phase Difference Induced from Charged Particles – A Josephson junction (JJ) consists of two superconductors with an insulating barrier in between. The electrons inside superconductors form Cooper pairs and become coherent states once the temperature is below the critical temperature. These two superconducting states on the different sides of the insulator, |$\ket{1}$ and |$\ket{2}$, can be described by macroscopic wave functions, $\Psi_{1,2}=\sqrt{n_{1,2}}e^{i\phi_{1,2}}$ Landau:1950lwq . A tunneling current occurs when there is a phase difference between the two states, $\Delta\phi\equiv\phi_{1}-\phi_{2}\neq 0$. This is the so-called Josephson effect Josephson:1962zz ; Josephson:1974uf . The Josephson current exhibits periodic dependence on phase, $I=I_{c}\sin(\Delta\phi)$ where $I_{c}$ is the constant critical current.
Figure 1: Schematic diagram illustrating the detection of moving charged particles by an RF-SQUID. When a charged particle passes by the insulator at a distance $d$, the resulting electric potential difference $\Delta V$ induces a phase change across the insulator, which in turn causes a magnetic flux and a variation in the Josephson inductance.
If the quantum states on each side of the insulator feel a voltage difference [^1] Voltage can also be replaced by any interaction potential involving electrons, such as gravitational potential Jain:1987zz ; Ummarino:2020loo ; Christodoulou:2023hyt ; Cheng:2024zde and the fifth-force potential Cheng:2024yrn . $\Delta V$, their phase difference can evolve over time according to the Schrodinger equation as $d\phi/dt=2e\Delta V$. The factor of 2 arises from the fact that a Cooper pair consists of two electrons. A charged particle passing by a JJ can induce a potential and phase difference across the junction. As shown in Fig. 1, an unknown charged particle $\chi$ flies past the junction at a distance of $d$ with a velocity $v_{\chi}$. Assuming the particle carries a charge $\epsilon e$, there exists a distance-dependent potential around it, with a magnitude of $V(r)=\epsilon e/(4\pi r)$. Due to the thickness $\delta$ of the insulator in JJ, there is a distance difference of $\delta$ between the superconductors on either side of the insulating layer and the charged particle. Since the electric field is shielded inside superconductors, a voltage difference only exists across the junction. This voltage is also a function of time because of the motion of charged particle. The accumulated phase difference is the integration over the entire time domain as,
$\Delta\phi=\frac{\epsilon e^{2}}{2\pi}\int\frac{1}{\sqrt{d^{2}+v_{\chi}^{2}t^{2}}}-\frac{1}{\sqrt{(d+\delta)^{2}+v_{\chi}^{2}t^{2}}}dt.$ (1)
The thickness of the insulating layer is typically $\delta=1$ nm, which is tiny compared to the overall size of a JJ ($\sim 10\,\mu$m) and the distance $d$ (which is $d\simeq 1\,$mm as explained in the section after the next). With $\delta\ll d$, the above integration can be calculated analytically,
$\Delta\phi=\frac{\epsilon e^{2}}{\pi}\frac{\delta}{v_{\chi}d}.$ (2)
Although the integration in Eq. (1) is performed from $t\rightarrow-\infty$ to $t\rightarrow+\infty$, the dominant contribution to the signal occurs when $\chi$ moves to a horizontal distance from the JJ that is less than $d$ (shown as the red dashed line in Fig. 1). During this distance, the travel time of $\chi$ is $\tau\simeq d/v_{\chi}$. Consequently, the flying charged particle induces a pulsed signal with frequency $f=1/\tau\simeq v_{\chi}/d$.
RF SQUID as Particle Detector – To detect the randomly occurring phase difference, we utilize the fact that a JJ behaves as a nonlinear inductance. The Josephson inductance, $L_{J}=L_{c}\sec(\phi)$, is a function of the phase difference $\phi=\phi_{0}+\Delta\phi$ across the junction, where $L_{c}=\Phi_{0}/(2\pi I_{c})$ and $\Phi_{0}\equiv\pi/e$. This expression holds true as long as the variation in the phase difference, $\Delta\phi$, is small around the equilibrium value $\phi_{0}$. Due to the perturbation from passing charged particles, a small phase difference $\Delta\phi$ develops across the junction, inducing a change in the Josephson inductance. The change in inductance can be detected by embedding the JJ into a superconducting loop with loop inductance $L_{S}$, forming an RF-SQUID (the blue circuit in Fig. 1). We choose $\lambda\equiv L_{S}/L_{C}<1$ to ensure a non-hysteretic RF-SQUID.
The general configuration and detectability of the chamber can be realized by a proposed microwave readout schema shown in Fig. 2. It was originally developed for large transition edge sensor (TES) arrays designed for cosmic microwave background observation and high-resolution X-ray/Gamma-ray spectrometers, where an RF-SQUID coupled to a TES pixel acts as a nonlinear flux-actuated inductor 2021ApPhL.118f2601D ; 6994798 ; 256711 ; 2017ApPhL.111f2601M ; HOLMES:2019ykt ; lens.org/027-935-516-419-894 ; SathyanarayanaRao:2020itw . In our proposal, charged particles directly interact with the JJ in the RF-SQUID, inducing a phase variation $\Delta\phi$ across the junction and altering its inductance. This results in a measurable resonance frequency shift of the quarter-wave resonators from $f_{0}$ to $f(\phi)$. In the small signal limit, the shift is Wegner:2021yrn ,
$f(\phi)=f_{0}-4f_{0}^{2}\left(C_{C}Z_{0}+\frac{L_{C}}{Z_{0}}-\frac{M_{C}^{2}}{L_{S}Z_{0}}\frac{\lambda\cos\phi}{1+\cos\phi}\right).$ (3)
Here, $f_{0}$ is the first mode resonance frequency of the unloaded resonator and $Z_{0}$ is its characteristic impedance. By measuring the resonance frequency shift, the phase variation $\Delta\phi$ induced by the charged particles can be deduced via the flux ramp modulation technique. Details are provided in Appendix A.
Figure 2: A schematic representation of a multiplexed readout module for the superconducting cloud chamber. The module comprises multiple microwave resonators capacitively coupled (with capacitance $C_{C}$) to a common microwave feedline. Each resonator is modulated by its own RF-SQUID, which serves as the fundamental particle detection unit. The RF-SQUID is inductively coupled (with mutual inductance $M_{C}$) to the load impedance $L_{C}$ of the quarter-wave resonator. The microwave signal transmitted through the feedline is amplified by a HEMT amplifier. Additionally, a common flux-ramp modulation line is applied to all RF-SQUIDs to linearize their responses.
Sensitivity – The resolution of the phase variation $\Delta\phi$ induced by the passing charged particles is determined by the flux noise of the readout module. Since the RF-SQUID used here is dissipationless and there is no input filter, not like in the readout setup for TES arrays, the primary noise contributions are from: (1) $1/f$ noise of the RF-SQUID, (2) Johnson noise of the microwave amplifier, (3) two-level system (TLS) noise of the quarter-wave resonator, and (4) other energy loss factors in the readout circuit. State-of-the-art readout modules have achieved a system noise of less than $1\mu\Phi_{0}/\rm{Hz}^{1/2}$ 2020ApPhL.117l2601N ; Malnou:2023wfo . Additionally, low-frequency noise is further suppressed by the flux ramp modulation technique.
In conclusion, within the frequency range from 1 Hz to MHz, the detector exhibits its optimal readout sensitivity $S\simeq 1\mu\Phi_{0}/\sqrt{1\,\text{Hz}}=2\times 10^{-21}\,\text{Wb}/\sqrt{1\,\text{Hz}}$ articlemates . This means that the detector is only sensitive to particles with velocities in the range of $v_{\chi}\subset(d/10^{-1}\,\text{s},d/10^{-6}\,\text{s}$). In our setup, the spacing $d=1\,$mm and thus the corresponding detectable particle velocity is quite small as, $v_{\chi}\subset(3\times 10^{-11},3\times 10^{-6})=(0.01\,\text{m}/\text{s},1000\,\text{m}/\text{s})$.
[…]
$\frac{2\pi\Delta\Phi}{\Phi_{0}}\simeq\Delta\phi\quad\rightarrow\quad\Delta\Phi\simeq\frac{\Phi_{0}\Delta\phi}{2\pi}.$ (4)
This magnetic flux can be understood from the following physical picture. The superconductors on both sides of the insulator are connected, so in the absence of an external field, their phases should be the same, $\phi_{2}=\phi_{1}$. However, due to perturbation caused by the passing charged particles, a phase difference and Josephson current are induced across the junction. This current generates a magnetic flux $\Delta\Phi$ through the self-inductance. The magnetic flux in the closed loop, in turn, produces an opposite phase difference, which offsets the phase difference induced by charged particles to maintain the aforementioned relationship [^2] Notice that the magnetic field of the low-speed particle is a secondary effect compared to its electric field. The magnetic flux is actually induced by the electric field of passing charged particles and the self-inductance of RF SQUID.. Therefore, the criterion for the signal to be observable is that it must be greater than the experimental sensitivity, $\Delta\Phi>S\times\sqrt{f}$.
Superconducting Cloud Chamber – Based on the particle-detector unit made of an RF SQUID in Fig. 1 and the readout schema in Fig. 2, we propose a superconducting cloud chamber consisting of a 3D array formed by these units for visualizing the trajectories of charged particles. As shown in Fig. 3, each cube represents an RF SQUID shown in Fig. 1, and the distance between each detection unit in all three directions is $D$. Typically, the size of an RF SQUID is $\mathcal{O}(10)\sim\mathcal{O}(100)\,\mu$m. We can design the spacing $D$ in the array to be $10$ times larger than the size of a single unit, such as $D\simeq 1\,$mm. In this way, if a particle passes through the cloud chamber, its distance $d$ from the closest SQUIDs will be fixed roughly as $d\simeq D=1\,$mm.
For example, to create a cube-shaped cloud chamber with a volume of 111\,1cm3, one needs $10\times 10\times 10=10^{3}$ SQUIDs in total. The design of the cloud chamber can help us easily identify the signals and trajectories of the particles passing through. Firstly, the detectors closest to the particle will measure the strongest signal. The remaining detectors will register progressively weaker signals as the distance increases. As depicted in Fig. 3, the black dashed line is the trajectory of a charged particle. The varying shades of red represent the strength of the signal, with darker red indicating a stronger signal.
Figure 3: The schematic diagram of the superconducting cloud chamber. Each cube represents a basic detection unit shown in Fig.1, forming a 3D array with a spacing of $D$ between adjacent units. When a charged particle passes through the cloud chamber, it excites pulsed signals in JJs surrounding its trajectory. By analyzing the signal strength and timing sequence, we can reconstruct the trajectory and velocity of the particle, as well as its coupling strength with electrons.
Furthermore, as previously mentioned, the pulsed signal is only generated when the particle enters the range $d\simeq D$ of the SQUID. Therefore, along the trajectory of the particle, the signals from the detector occur almost sequentially in time. Therefore, we can reconstruct the trajectory and velocity of the particle by analyzing the signal strength and time sequence of all detectors  [^3] In principle, we can increase the size of the chamber to enhance its ability to detect trajectories. However, current technology is not able to connect a large number of SQUIDs in a single circuit and effectively read them out. We can divide the setup into many smaller chambers and arrange them in regular pattern to achieve the same effect. The maximum number of SQUIDs in a readout module is discussed in Appendix B. .
The superconducting cloud chamber can also effectively eliminate the majority of environmental backgrounds. Firstly, random system backgrounds only sporadically generate signals and do not present particle trajectories. Secondly, since the detector is sensitive only to low-speed flying charged particles, signals generated by high-energy SM particles like cosmic rays cannot be detected, thus not forming background noise [^4] Although cosmic muons may be decelerated by Earth shielding, they quickly decay to electrons and the speed of electrons is too high.. Lastly, once clear particle trajectories are identified in the chamber, confirming the presence of flying particles, we can also determine the charge of particle by analyzing the signal intensities. In the case of millicharged DM detection, the charges of SM charged particles in the background are all integers, making them easily distinguishable.
Application to Millicharged DM – As mentioned in the introduction, one application of this chamber is the detection of millicharged DM. Previous research Neufeld:2018slx ; Pospelov:2020ktu ; Leane:2022hkk ; Berlin:2023zpn has shown that long-range and strongly interacting DM can reach thermal equilibrium through multiple scatterings with the Earth’s environment. The millicharged DM is assumed to directly couples only to massless SM photons. Once its electric charge $\epsilon$ satisfies the condition Pospelov:2020ktu , $\epsilon\gtrsim 3\times 10^{-6}\sqrt{m_{\chi}/\rm{GeV}}$, it can reach thermal equilibrium at depths of 111\,1km underground (the black solid line in Fig. 4). As a result, the kinetic energy of DM equals the Earth’s temperature, approximately $T_{\chi}=0.2\,$eV. The velocity of DM relates to its mass as $v_{\chi}=\sqrt{2T_{\chi}/m_{\chi}}$.
Figure 4: The projected sensitivity of superconducting cloud chamber on millicharged DM, shown as the different colored shaded region depending on if the DM is thermalized or not. The constraint from collider is the grey shaded region, while the constraints from ion trap experiments are blue solid and green dashed lines depending on the DM fraction.
The detectable velocity range of a charged particle in superconducting cloud chamber, $(3\times 10^{-11},3\times 10^{-6})$, corresponds to a detectable range of the millicharged DM masses, $m_{\chi}\sim(1,10^{10})\,$GeV. The minimum magnetic flux that the detector can measure depends on the signal frequency, or equivalently, the DM velocity,
$\delta\Phi=\frac{1\mu\Phi_{0}}{\sqrt{\text{Hz}}}\sqrt{\frac{v_{\chi}}{1\,\text{mm}}}=0.5\times\Phi_{0}\sqrt{v_{\chi}}.$ (5)
Compare it with the signal Eq. (4), one can get the sensitivity of our cloud chamber to the phase difference $\Delta\phi\geq\pi\sqrt{v_{\chi}}$ and then the millicharge,
$\epsilon\geq 10^{8}\times v_{\chi}^{3/2}=17.8\times\left(\frac{0.4\,\rm{GeV}}{m_{\chi}}\right)^{3/4},$ (6)
which is the region above the red solid line in Fig. 4. Since this limit only applies to the DM thermalized with the Earth environment, its overlap (red shaded region) with region above black line is the sensitivity of the present superconducing cloud chamber to millicharged DM.
[…]
For the DM that can not fully reached thermalization at a distance of 111\,1km inside the Earth (black dashed line), a small fraction of them still have lower velocities due to distribution effects. Since the chamber has very low requirements for the DM number density, this parameter space (orange shaded region) might also be detectable.
Due to its extremely low kinetic energy, the millicharged DM thermalized in Earth cannot surpass the threshold for traditional DM direct detections Budker:2021quh ; ArguellesDelgado:2021lek ; PandaX:2023toi ; Iles:2024zka , necessitating novel experimental methods for its detection, such as electrostatic accelerator Pospelov:2020ktu , oil-drop experiment Kim:2007zzs , levitation experiments with microspheres Moore:2014yba ; Afek:2020lek , ion traps Budker:2021quh , and so on Chen:2022abz ; CUORE:2024rbd ; Ema:2024oce . For DM fraction of $f_{\chi}=10^{-6}$ and $f_{\chi}=10^{-12}$, the most stringent constraints from ion-trap experiments are shown as the blue solid and green dashed lines Budker:2021quh . Because of the decrease in energy transfer efficiency and number density, this limit is not sensitive to high-mass and small-fraction parameter ranges. We also demonstrate collider search and beam dump experiment constraints as the gray region 1995ZPhyC..67..203A ; Prinz:1998ua ; Magill:2018tbb ; Berlin:2018bsc ; Kelly:2018brz ; Harnik:2019zee ; ArgoNeuT:2019ckq ; Ball:2020dnx ; Marocco:2020dqu . It can be observed that the superconducting cloud chamber has a unique advantage in the high-mass parameter range, as higher-mass millichagred DM particles have slower velocities, aligning perfectly with the sensitivity range of the detector.
Discussions and Summary – Particle tracking detectors play a crucial role in the discovery of new particles, greatly advancing the fields of particle physics. Many fundamental charged particles such as the positron and muon were discovered in Wilson cloud chamber Grupen_Shwartz_2008 . Afterwards, the bubble chamber made a series of significant discoveries Glaser:1952zz including weak neutral currents. Cherenkov detectors can also be used to detect neutral high-speed particles, such as high-energy neutrinos Bolotovskii_2009 . Recently, new tracking detectors such as wire chambers, spark chambers, drift chambers, and solid-state nuclear track detectors have been developed for a wider range of physics explorations Sauli:1977mt ; GUO2012233 . DM is typically considered as an unknown particle. Its detection can also be aided by various bubble chambers PICASSO:2012ngj ; COUPP:2012jrk ; PICO:2015amc ; PICO:2023uff .
The detection methods based on scattering and ionization used by the detectors above all have certain energy thresholds. Ionizing an atom typically requires $\sim 10\,$eV of energy. This requires the kinetic energy of the incident particles to be greater than this threshold. Consequently, they can not detect the millicharged DM thermalized with the Earth’s environment and even the SM charged particles with velocity lower than $10^{-4}$.
In this paper, we propose a cloud chamber made of JJs. Due to the absence of resistance in superconductors, the energy threshold required to excite superconducting current signals is extremely low. Considering the background of circuit readout, superconducting cloud chamber can detect charged or millicharged particles traveling at speeds of $v_{\chi}\sim(10^{-11},10^{-6})$. Even if it constitutes only $f_{\chi}\gtrsim 10^{-12}$ of DM, the present superconducting cloud chamber is sensitive to millicarged DM with masses of $(10^{5},10^{10})\,$GeV. Furthermore, this chamber can also be used to track the movement trajectories of charged particles in space or specific laboratories, opening up new avenues for the detection of low-speed charged particles.
Acknowledgements
The authors thank Tie-Sheng Yang for improving the Fig.$1\sim 3$. They also thank Hong Ding, Shao-Feng Ge, Shigeki Matsumoto, and Ning Zhou for fruitful discussions. B. G. is supported by the Innovation Program for Quantum Science and Technology (No.2021ZD0302700). J. S. is supported by the National Natural Science Foundation of China (Nos. 12375101, 12425506, 12090060, 12090064) and the SJTU Double First Class start-up fund WF220442604. T. T. Y. is supported by the Natural Science Foundation of China (NSFC) under Grant No. 12175134, MEXT KAKENHI Grants No. 24H02244, and World Premier International Research Center Initiative (WPI Initiative), MEXT, Japan.
Appendix A Appendix A – Detection of the Phase Variation
The RF-SQUID exhibits a non-linear response to charged particles. To linearize this response, a flux-ramp modulation technique is employed. A sawtooth signal with frequency $f_{\rm ramp}$ and amplitude $n\Phi_{0}$ (where n is an integer) is applied to the common flux-ramping line. This induces periodic oscillations in the SQUID response. If the slew rate of the applied ramp exceeds that of any input signal, variations in the input signal appear as phase offsets during the ramp period, producing a phase shift in the SQUID response relative to its free oscillation state (i.e., without the interaction of charged particles). This phase shift can be calculated from a Fourier measurement of the SQUID response,
$\phi=\arctan\left(\frac{-\sum\theta(t)\sin(2\pi f_{\pi}t)}{-\sum\theta(t)\cos(2\pi f_{\pi}t)}\right).$ (7)
Here, $\theta(t)$ represents the discrete sampling of the SQUID response, and $f_{m}=n\times f_{\rm ramp}$ is the modulation frequency. Typically, $f_{m}$ is required to be at least $10$ times the maximum signal frequency.
Appendix B Appendix B – Multiplexing Factor
The number of RF SQUIDs in a readout module, which we define as the multiplexing factor $M$, can be estimated using as follows. Let $S_{f}$ be the frequency spacing between resonator tones, and $f_{\rm ADC}$ be the ADC bandwidth, the multiplexing factor is then defined as $M\equiv f_{\rm ADC}/S_{f}$. Several constraints apply to $S_{f}$ HOLMES:2019ykt : (1) $S_{f}$ must be significantly larger than the resonator bandwidth, $S_{f}>g_{f}\times\Delta f_{\rm BW}$, where $\Delta f_{\rm BW}$ is the resonantor bandwidth and $g_{f}$ is a guard factor; (2) The Nyquist-Sannon sampling theorem requires $\Delta f_{\rm BW}>2f_{m}$; (3) The ramp frequency must exceed the maximum signal frequency, $f_{\rm ramp}>R_{d}/\tau$, where $\tau$ is the rise time of the charged particle (DM) signal and $R_{d}$ is a signal distortion factor. In conclusion, the multiplexing factor per ADC board is,
$M=\frac{f_{\rm ADC}\cdot\tau}{2n\cdot g_{f}\cdot R_{d}}.$ (8)
Assuming the typical values of $f_{\rm ADC}=1\,$GHz, $n=5$, $g_{f}=5$, and $R_{d}=5$, if a signal rise time is $10\,\mu$s, the multiplexing factor per ADC board is estimated to be $M=40$. If the frequency of signal becomes lower and the rise time is extended to be 111\,1ms, the multiplexing factor increases to be $M=4000$. This multiplexing capability, which depends on the frequency of signals, has been demonstrated in TES-based CMB detector arrays Henderson:2018jlc ; Jones:2024yss .

### v2

@@
 By measuring the quantum phase difference, this device can detect charged particles with extremely low kinetic energy, providing a new method for detecting slow-moving particles.
-It can also be utilized to detect millicharged dark matter particles thermalized with the Earth’s environment within the mass range of $10^{5}\sim 10^{10}\,$eV.
+It can also be utilized to detect millicharged dark matter particles thermalized with the Earth’s environment within the mass range of $10^{4}\sim 10^{10}\,$eV.
 Introduction – Millicharged particles can be an interesting dark matter (DM) candidate Holdom:1985ag ; Feldman:2007wj ; McDermott:2010pa ; Izaguirre:2015eya ; Agrawal:2016quu because millicharge naturally ensures the stability of particle.
@@
 Consequently, the kinetic energy of DM becomes as tiny as the Earth’s temperature ($\sim 0.2\,$eV), which prevents it from surpassing the energy threshold of DM direct detection experiments and thus evading experimental constraints Pospelov:2020ktu ; Berlin:2023zpn .
-Additionally, such a strong-interacting millicharged DM is thermally produced in the early Universe with a small abundance Aboubrahim:2021ohe .
+Additionally, these strongly-interacting millicharged DM particles can be thermally produced in the early Universe with a small abundance Aboubrahim:2021ohe .
 The cosmological observations are not sensitive to the DM with a small fraction Dubovsky:2003yn ; dePutter:2018xte ; Kovetz:2018zan ; Buen-Abad:2021mvc .
@@
 Moreover, current particle detectors are unable to directly observe extremely low-velocity particles, even those SM charged particles, without acceleration.
-In this paper, we design a novel superconducting cloud chamber made of RF SQUIDs (radio frequency superconducting quantum interference device 34120 ) to target low-speed charged particles.
+In this paper, we design a novel superconducting cloud chamber made of RF SQUIDs (radio frequency superconducting quantum interference devices 34120 ) to target low-speed charged particles.
 We use natural units with $c=\hbar=1$ throughout the text.
@@
 This voltage is also a function of time because of the motion of charged particle.
-The accumulated phase difference is the integration over the entire time domain as,
+When the distance $d$ is much larger than the planar dimensions of the JJ, the JJ can be taken as one-dimensional and the accumulated phase difference is the integration over the entire time domain as,
 $\Delta\phi=\frac{\epsilon e^{2}}{2\pi}\int\frac{1}{\sqrt{d^{2}+v_{\chi}^{2}t^{2}}}-\frac{1}{\sqrt{(d+\delta)^{2}+v_{\chi}^{2}t^{2}}}dt.$ (1)
@@
 During this distance, the travel time of $\chi$ is $\tau\simeq d/v_{\chi}$.
-Consequently, the flying charged particle induces a pulsed signal with frequency $f=1/\tau\simeq v_{\chi}/d$.
+Consequently, the flying charged particle induces a pulsed signal with frequency $f_{\text{signal}}=1/\tau\simeq v_{\chi}/d$.
+[…]
+$\frac{2\pi\Delta\Phi}{\Phi_{0}}\simeq\Delta\phi\quad\rightarrow\quad\Delta\Phi\simeq\frac{\Phi_{0}\Delta\phi}{2\pi}.$ (3)
 RF SQUID as Particle Detector – To detect the randomly occurring phase difference, we utilize the fact that a JJ behaves as a nonlinear inductance.
@@
 It was originally developed for large transition edge sensor (TES) arrays designed for cosmic microwave background observation and high-resolution X-ray/Gamma-ray spectrometers, where an RF-SQUID coupled to a TES pixel acts as a nonlinear flux-actuated inductor 2021ApPhL.118f2601D ; 6994798 ; 256711 ; 2017ApPhL.111f2601M ; HOLMES:2019ykt ; lens.org/027-935-516-419-894 ; SathyanarayanaRao:2020itw .
+The RF-SQUID is coupled to the load inductor of a quarter-wave resonator, such that the input signal from the TES modulates the resonance frequency of resonator.
 In our proposal, charged particles directly interact with the JJ in the RF-SQUID, inducing a phase variation $\Delta\phi$ across the junction and altering its inductance.
-This results in a measurable resonance frequency shift of the quarter-wave resonators from $f_{0}$ to $f(\phi)$.
+This results in a measurable resonance frequency shift of the quarter-wave resonators from $f_{0}$ to $f(\Delta\phi)$.
 In the small signal limit, the shift is Wegner:2021yrn ,
-$f(\phi)=f_{0}-4f_{0}^{2}\left(C_{C}Z_{0}+\frac{L_{C}}{Z_{0}}-\frac{M_{C}^{2}}{L_{S}Z_{0}}\frac{\lambda\cos\phi}{1+\cos\phi}\right).$ (3)
+$f(\Delta\phi)=f_{0}-4f_{0}^{2}\left(C_{C}Z_{0}+\frac{L_{C}}{Z_{0}}-\frac{M_{C}^{2}}{L_{S}Z_{0}}\frac{\lambda\cos\Delta\phi}{1+\cos\Delta\phi}\right).$ (4)
 Here, $f_{0}$ is the first mode resonance frequency of the unloaded resonator and $Z_{0}$ is its characteristic impedance.
-By measuring the resonance frequency shift, the phase variation $\Delta\phi$ induced by the charged particles can be deduced via the flux ramp modulation technique.
-Details are provided in Appendix A.
+By measuring the resonance frequency shift, the phase variation $\Delta\phi$ induced by the charged particles can be inferred.
+In practice, a specialized readout technique known as the flux ramp has been developed.
+The core idea is to apply a sawtooth waveform to the RF-SQUID via a flux-ramping line.
+The ramp frequency of the sawtooth waveform, $f_{\text{ramp}}$, is chosen to be at least twice the frequency of the input signal.
+The amplitude of the waveform is set such that it induces an integer number of flux quanta in the RF-SQUID.
+This waveform periodically modulates the response of the RF-SQUID at a modulation frequency $f_{\text{mod}}=nf_{\text{ramp}}$, and consequently modulates the resonance frequency of the quarter-wave resonator.
+Since $f_{\text{mod}}$ is much higher than the signal frequency, the influence of the input signal appears as a phase shift in the modulated SQUID response.
+It is from this phase shift that the signal of a potential charged particle can be extracted.
+Further details are provided in Appendix A.
 Figure 2: A schematic representation of a multiplexed readout module for the superconducting cloud chamber.
@@
 Since the RF-SQUID used here is dissipationless and there is no input filter, not like in the readout setup for TES arrays, the primary noise contributions are from: (1) $1/f$ noise of the RF-SQUID, (2) Johnson noise of the microwave amplifier, (3) two-level system (TLS) noise of the quarter-wave resonator, and (4) other energy loss factors in the readout circuit.
-State-of-the-art readout modules have achieved a system noise of less than $1\mu\Phi_{0}/\rm{Hz}^{1/2}$ 2020ApPhL.117l2601N ; Malnou:2023wfo .
+State-of-the-art readout modules have achieved a system white noise of less than $1\mu\Phi_{0}/\rm{Hz}^{1/2}$ 2020ApPhL.117l2601N ; Malnou:2023wfo .
 Additionally, low-frequency noise is further suppressed by the flux ramp modulation technique.
-In conclusion, within the frequency range from 1 Hz to MHz, the detector exhibits its optimal readout sensitivity $S\simeq 1\mu\Phi_{0}/\sqrt{1\,\text{Hz}}=2\times 10^{-21}\,\text{Wb}/\sqrt{1\,\text{Hz}}$ articlemates .
-This means that the detector is only sensitive to particles with velocities in the range of $v_{\chi}\subset(d/10^{-1}\,\text{s},d/10^{-6}\,\text{s}$).
+As explained above, the signal frequency is limited by the ramp frequency $f_{\text{ramp}}$, which cannot be set arbitrarily high because the resulting modulation frequency $f_{\text{mod}}$ is constrained by the bandwidth of the quarter-wave resonator.
+We take the benchmark value of the ramp frequency as $f_{\text{ramp}}=1\,$MHz.
+On the other hand, readout noise increases at low frequencies due to $1/f$ noise.
+The corner frequency, $f_{\text{knot}}$, where $1/f$ noise equals the white noise level, can be as low as $f_{\text{knot}}=10\,$Hz.
+Therefore, the detectable signal-frequency-range is effectively limited to the band between $f_{\text{knot}}$ and $f_{\text{ramp}}$ as $(10,10^{6})\,$Hz.
+As a result, the entire readout chain can be band-limited to this frequency range, and noise components beyond $f_{\text{ramp}}$ will be discarded.
+Within this frequency range, the total readout noise can be estimated using,
+$\Phi_{\text{n}}\equiv 1\mu\Phi_{0}/\sqrt{\text{Hz}}\times\sqrt{f_{\text{ramp}}}=10^{-3}\times\Phi_{0}.$ (5)
+Based on the frequency relationship of the signal, $f_{\text{signal}}=v_{\chi}/d$, the aforementioned signal-frequency-range can be further converted into the detectable velocity range of the particles, $d/10^{-1}\,\text{s}<v_{\chi}<d/10^{-6}\,\text{s}$.
 In our setup, the spacing $d=1\,$mm and thus the corresponding detectable particle velocity is quite small as, $v_{\chi}\subset(3\times 10^{-11},3\times 10^{-6})=(0.01\,\text{m}/\text{s},1000\,\text{m}/\text{s})$.
-[…]
-$\frac{2\pi\Delta\Phi}{\Phi_{0}}\simeq\Delta\phi\quad\rightarrow\quad\Delta\Phi\simeq\frac{\Phi_{0}\Delta\phi}{2\pi}.$ (4)
-This magnetic flux can be understood from the following physical picture.
-The superconductors on both sides of the insulator are connected, so in the absence of an external field, their phases should be the same, $\phi_{2}=\phi_{1}$.
-However, due to perturbation caused by the passing charged particles, a phase difference and Josephson current are induced across the junction.
-This current generates a magnetic flux $\Delta\Phi$ through the self-inductance.
-The magnetic flux in the closed loop, in turn, produces an opposite phase difference, which offsets the phase difference induced by charged particles to maintain the aforementioned relationship [^2] Notice that the magnetic field of the low-speed particle is a secondary effect compared to its electric field.
-The magnetic flux is actually induced by the electric field of passing charged particles and the self-inductance of RF SQUID..
-Therefore, the criterion for the signal to be observable is that it must be greater than the experimental sensitivity, $\Delta\Phi>S\times\sqrt{f}$.
-Superconducting Cloud Chamber – Based on the particle-detector unit made of an RF SQUID in Fig.
-1 and the readout schema in Fig.
+Since the noise $\Phi_{n}$ in Eq.
+(5) is expressed in terms of magnetic flux, the detectability of a signal should also be assessed by comparing it to the magnetic flux signal $\Delta\Phi$ generated by the charged particle in Eq.
+(3).
+Taking $f_{\text{ramp}}=1\,$MHz, the signal-to-noise ratio (SNR) can therefore be estimated as,
+$\text{SNR}\equiv\frac{\Delta\Phi}{\Phi_{n}}=\frac{\Delta\Phi}{10^{-3}\times\Phi_{0}}.$ (6)
+Conservatively, we consider the criterion for detectable signal is that $\text{SNR}\geq 10$ Qin:2023urf .
+Superconducting Cloud Chamber – Based on the particle-detector unit made of an RF SQUID and the readout schema in Fig.
 2, we propose a superconducting cloud chamber consisting of a 3D array formed by these units for visualizing the trajectories of charged particles.
@@
 1, and the distance between each detection unit in all three directions is $D$.
-Typically, the size of an RF SQUID is $\mathcal{O}(10)\sim\mathcal{O}(100)\,\mu$m.
-We can design the spacing $D$ in the array to be $10$ times larger than the size of a single unit, such as $D\simeq 1\,$mm.
-In this way, if a particle passes through the cloud chamber, its distance $d$ from the closest SQUIDs will be fixed roughly as $d\simeq D=1\,$mm.
-For example, to create a cube-shaped cloud chamber with a volume of 111\,1cm3, one needs $10\times 10\times 10=10^{3}$ SQUIDs in total.
+The geometrical size of a non-hysteresis RF-SQUID with $L_{S}<L_{J}$ must be small, on the order of $\mathcal{O}(10)\,\mu$m, and the dimension of JJ is $a\simeq 3\,\mu$m.
+We can design the spacing $D$ in the array to be tens of times larger than the size of a single unit, such as $D\simeq 3\,$mm.
+This means each JJ is surrounded by an exclusion volume $(3\,\text{mm})^{3}$, making the probability of a particle passing in close proximity to a JJ extremely low.
+For instance, the probability for a DM particle to approach a JJ within a distance $10a$ is $(30\,\mu\text{m})^{3}/(3\,\text{mm})^{3}=10^{-6}$.
+In this way, if a particle passes through the cloud chamber, its distance $d$ from the closest SQUIDs will be fixed roughly as $d\sim 1\,$mm $\lesssim D$.
+Therefore, the assumption of neglecting the planar size of JJ in Eq.
+(1) is well justified.
+For example, to create a cube-shaped cloud chamber with a volume of (333\,3cm)3, one needs $10\times 10\times 10=10^{3}$ SQUIDs in total.
 The design of the cloud chamber can help us easily identify the signals and trajectories of the particles passing through.
@@
 By analyzing the signal strength and timing sequence, we can reconstruct the trajectory and velocity of the particle, as well as its coupling strength with electrons.
-Furthermore, as previously mentioned, the pulsed signal is only generated when the particle enters the range $d\simeq D$ of the SQUID.
+Furthermore, as previously mentioned, the pulsed signal is only generated when the particle enters the range $d\simeq 1\,$mm of the SQUID.
 Therefore, along the trajectory of the particle, the signals from the detector occur almost sequentially in time.
@@
 In the case of millicharged DM detection, the charges of SM charged particles in the background are all integers, making them easily distinguishable.
+A detailed discussion on determining the velocity and charge of particles using the chamber can be found in Appendix.C.
 Application to Millicharged DM – As mentioned in the introduction, one application of this chamber is the detection of millicharged DM.
@@
 The velocity of DM relates to its mass as $v_{\chi}=\sqrt{2T_{\chi}/m_{\chi}}$.
-Figure 4: The projected sensitivity of superconducting cloud chamber on millicharged DM, shown as the different colored shaded region depending on if the DM is thermalized or not.
+Figure 4: The projected sensitivity of superconducting cloud chamber on millicharged DM assuming the $\text{SNR}=10$ (SNR10), shown as the different colored shaded region depending on if the DM is thermalized or not.
 The constraint from collider is the grey shaded region, while the constraints from ion trap experiments are blue solid and green dashed lines depending on the DM fraction.
+The sensitivity at $\text{SNR}=1$ is also displayed as a red dashed line.
 The detectable velocity range of a charged particle in superconducting cloud chamber, $(3\times 10^{-11},3\times 10^{-6})$, corresponds to a detectable range of the millicharged DM masses, $m_{\chi}\sim(1,10^{10})\,$GeV.
-The minimum magnetic flux that the detector can measure depends on the signal frequency, or equivalently, the DM velocity,
-$\delta\Phi=\frac{1\mu\Phi_{0}}{\sqrt{\text{Hz}}}\sqrt{\frac{v_{\chi}}{1\,\text{mm}}}=0.5\times\Phi_{0}\sqrt{v_{\chi}}.$ (5)
-Compare it with the signal Eq. (4), one can get the sensitivity of our cloud chamber to the phase difference $\Delta\phi\geq\pi\sqrt{v_{\chi}}$ and then the millicharge,
-$\epsilon\geq 10^{8}\times v_{\chi}^{3/2}=17.8\times\left(\frac{0.4\,\rm{GeV}}{m_{\chi}}\right)^{3/4},$ (6)
+Assuming the SNR to be $10$, the minimum magnetic flux that the detector can measure is, $\Delta\Phi\geq 10^{-2}\times\Phi_{0}$, according to Eq.
+(6).
+Compare it with the signal Eq. (3), one can get the sensitivity of our cloud chamber to the phase difference $\Delta\phi\geq\pi/50$ and then the millicharge,
+$\epsilon\geq 10^{9}\times v_{\chi}^{3/2}=39.4\times\left(\frac{\rm{GeV}}{m_{\chi}}\right)^{1/2},$ (7)
 which is the region above the red solid line in Fig.
@@
 We also demonstrate collider search and beam dump experiment constraints as the gray region 1995ZPhyC..67..203A ; Prinz:1998ua ; Magill:2018tbb ; Berlin:2018bsc ; Kelly:2018brz ; Harnik:2019zee ; ArgoNeuT:2019ckq ; Ball:2020dnx ; Marocco:2020dqu .
-It can be observed that the superconducting cloud chamber has a unique advantage in the high-mass parameter range, as higher-mass millichagred DM particles have slower velocities, aligning perfectly with the sensitivity range of the detector.
+It can be observed that the superconducting cloud chamber has a unique advantage in the high-mass parameter range, as higher-mass millicharged DM particles have slower velocities, aligning perfectly with the sensitivity range of the detector.
 Discussions and Summary – Particle tracking detectors play a crucial role in the discovery of new particles, greatly advancing the fields of particle physics.
@@
 Considering the background of circuit readout, superconducting cloud chamber can detect charged or millicharged particles traveling at speeds of $v_{\chi}\sim(10^{-11},10^{-6})$.
-Even if it constitutes only $f_{\chi}\gtrsim 10^{-12}$ of DM, the present superconducting cloud chamber is sensitive to millicarged DM with masses of $(10^{5},10^{10})\,$GeV.
+Even if it constitutes only $f_{\chi}\gtrsim 10^{-12}$ of DM, the present superconducting cloud chamber is sensitive to millicarged DM with masses of $(10^{4},10^{10})\,$GeV.
 Furthermore, this chamber can also be used to track the movement trajectories of charged particles in space or specific laboratories, opening up new avenues for the detection of low-speed charged particles.
@@
 This phase shift can be calculated from a Fourier measurement of the SQUID response,
-$\phi=\arctan\left(\frac{-\sum\theta(t)\sin(2\pi f_{\pi}t)}{-\sum\theta(t)\cos(2\pi f_{\pi}t)}\right).$ (7)
-Here, $\theta(t)$ represents the discrete sampling of the SQUID response, and $f_{m}=n\times f_{\rm ramp}$ is the modulation frequency.
-Typically, $f_{m}$ is required to be at least $10$ times the maximum signal frequency.
+$\phi=\arctan\left(\frac{-\sum\theta(t)\sin(2\pi f_{\text{mod}}t)}{-\sum\theta(t)\cos(2\pi f_{\text{mod}}t)}\right).$ (8)
+Here, $\theta(t)$ represents the discrete sampling of the SQUID response, and $f_{\text{mod}}=n\times f_{\rm ramp}$ is the modulation frequency.
+Typically, $f_{\text{mod}}$ is required to be at least $10$ times the maximum signal frequency.
 Appendix B Appendix B – Multiplexing Factor
@@
 Let $S_{f}$ be the frequency spacing between resonator tones, and $f_{\rm ADC}$ be the ADC bandwidth, the multiplexing factor is then defined as $M\equiv f_{\rm ADC}/S_{f}$.
-Several constraints apply to $S_{f}$ HOLMES:2019ykt : (1) $S_{f}$ must be significantly larger than the resonator bandwidth, $S_{f}>g_{f}\times\Delta f_{\rm BW}$, where $\Delta f_{\rm BW}$ is the resonantor bandwidth and $g_{f}$ is a guard factor; (2) The Nyquist-Sannon sampling theorem requires $\Delta f_{\rm BW}>2f_{m}$; (3) The ramp frequency must exceed the maximum signal frequency, $f_{\rm ramp}>R_{d}/\tau$, where $\tau$ is the rise time of the charged particle (DM) signal and $R_{d}$ is a signal distortion factor.
+Several constraints apply to $S_{f}$ HOLMES:2019ykt : (1) $S_{f}$ must be significantly larger than the resonator bandwidth, $S_{f}>g_{f}\times\Delta f_{\rm BW}$, where $\Delta f_{\rm BW}$ is the resonantor bandwidth and $g_{f}$ is a guard factor; (2) The Nyquist-Sannon sampling theorem requires $\Delta f_{\rm BW}>2f_{\text{mod}}$; (3) The ramp frequency must exceed the maximum signal frequency, $f_{\rm ramp}>R_{d}/\tau$, where $\tau$ is the rise time of the charged particle (DM) signal and $R_{d}$ is a signal distortion factor.
 In conclusion, the multiplexing factor per ADC board is,
@@
 This multiplexing capability, which depends on the frequency of signals, has been demonstrated in TES-based CMB detector arrays Henderson:2018jlc ; Jones:2024yss .
+Appendix C Appendix C – Determination of Particle Nature
+In the main text, we mentioned that by analyzing the signal amplitudes and timing sequences from multiple RF-SQUIDs in the superconducting cloud chamber, we can reconstruct a particle’s trajectory, charge, and velocity.
+In this appendix, we demonstrate how to extract this information by analyzing signals from any four arbitrarily selected JJs within a localized region of the chamber.
+Figure 5: A schematic illustration of a particle passing through the superconducting cloud chamber (local view).
+As shown in Fig.
+5, the charged particle flies by the detector $1\sim 4$ with distance $d_{i}(i=1\sim 4)$.
+Correspondingly, the detectors see the phase difference $\phi_{i}(i=1\sim 4)$ and the signal peak arises at the moment $t_{i}(i=1\sim 4)$.
+According to Eq.
+(2), we can list the following four equations,
+$\phi_{i}=\frac{\epsilon e^{2}}{\pi}\frac{\delta}{v_{\chi}d_{i}}.$ (10)
+However, the particle charge $\epsilon$ and velocity $v_{\chi}$ are degenerate information that cannot be directly determined from the equations above.
+At this stage, by introducing the assumption that the particle’s trajectory forms an angle $\theta$ with the detector array, we can establish the relationship between the measured distances,
+$d_{3}=d_{1}+\frac{D}{\cos\theta}\,,\quad d_{4}=d_{2}+\frac{D}{\cos\theta}.$ (11)
+We can also establish the relationship between the signal timing and the particle’s velocity,
+$t_{2}-t_{1}=\frac{D}{\cos\theta v_{\chi}}.$ (12)
+We have now formulated a system of seven independent equations, enabling the determination of all seven unknowns, $\epsilon$, $v_{\chi}$, $\theta$, and $d_{i}(i=1\sim 4)$.
+From them, we can figure out the charge and velocity of the flying particle.
+To see the time difference, the timing resolution between different detectors must be better than, $\delta t<t_{2}-t_{1}\simeq D/v_{\chi}$.
+Since $D\sim 1\,$mm and $v_{\chi}\subset(0.01,1000)\,$m/s, the time resolution should be better than $\delta t<10^{-5}\,$s.
+This condition is easily achieved by the current timing technology.